Cal Farms reported supplies expense of $1,700,000 this year. The supplies account decreased by $190,000 during the year to an ending balance of $340,000. What was the cost of supplies the Cal Farms purchased during the year?

Multiple Choice

  • $1,510,000.

  • $1,360,000.

  • $1,890,000.

  • $2,040,000.
Answer #1

Beginning supplies balance=(190,000+340,000)=$530,000

Ending supplies balance=Beginning supplies balance+Cost of supplies purchased-Supplies expense

340,000=530,000+Cost of supplies purchased-1,700,000

Cost of supplies purchased=340,000+1,700,000-530,000

=$1,510,000.